OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. -(1/16)x^2 = y

OpenStudy (misssunshinexxoxo):

x^2=−16y

OpenStudy (misssunshinexxoxo):

Plug in (y - k) = a (x - h)^2 and gives solution :)

OpenStudy (perl):

You can use the equation $$ \Large y= \frac{1}{4p} (x-h)^2 +k $$The distance from the origin \( (h,k) \) to the focus is \( |p | \)